Evaluate the following infinite series. so (-1)+12n-1 2n=0 n! 2e-2 e-2 0 -2e-2 0-5e-2 O none of the above or below
Answer #1

we are given series as

\sum_{n=0}^{\infty}\frac{(-1)^{n+1}2^{n-1}}{n!}

we can use series formula

\sum_{n=0}^{\infty}\frac{x^n}{n!}=e^x

we can write as

\sum_{n=0}^{\infty}\frac{(-1)^{n+1}2^{n-1}}{n!}=\sum_{n=0}^{\infty}\frac{(-1)(-1)^{n}2^{n}}{2(n!)}

\sum_{n=0}^{\infty}\frac{(-1)^{n+1}2^{n-1}}{n!}=-\frac{1}{2}\sum_{n=0}^{\infty}\frac{(-1)^{n}2^{n}}{(n!)}

\sum_{n=0}^{\infty}\frac{(-1)^{n+1}2^{n-1}}{n!}=-\frac{1}{2}\sum_{n=0}^{\infty}\frac{(-2)^{n}}{(n!)}

we can plug x=-2 in series formula

and we get

\sum_{n=0}^{\infty}\frac{(-1)^{n+1}2^{n-1}}{n!}=-\frac{1}{2}e^{-2}...........Answer